Circular No. 7458

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cx IN NGC 524 C. Yu, M. Modjaz, and W. D. Li, University of California at Berkeley,

on behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on July 17.5 (mag ∼ 14.5) and 18.4 UT (mag ∼ 14.3). The new object is located at α = 1^{h}24^{m}46^{s}.15, δ = +9^{o}30′30′′.9 (equinox 2000.0), which is 23′′.0 west and 109′′.3 south of the nucleus of NGC 524. A KAIT image of the same field taken on 1998 Aug. 31.4 showed nothing at the position of SN 2000cx (limiting mag ∼ 19.5).

NOVA IN NGC 185 Modjaz and Li also report the discovery with KAIT of an apparent nova

on unfiltered images taken on July 13.5 (mag ∼ 18.7) and 14.5 UT (mag ∼ 18.2). The new object is located at α = 0^{h}39^{m}07^{s}.02, δ = +48^{o}20′34′′.4 (equinox 2000.0), which is ∼ 93′′ east and 20′′ north of the diffuse nucleus of NGC 185. A KAIT image of the same field on July 11.5 (limiting mag ∼ 19.0) showed nothing at the position of the nova.

COMETS C/1999 G3, C/1999 G4, C/1999 H6 (SOHO) D. Hammer reports measurements of three more Kreutz sungrazing

comets from SOHO LASCO C2 data (cf., e.g., IAUC 7452, 7454), all found by M. Oates from archival data on the SOHO website. The discovery observations are given below, while MPEC 2000-O06 contains the fully reduced observations and orbital elements (by B. G. Marsden).

1999 UT  α_{2000}  δ_{2000}  Comet
Apr. 2.021 0^{h}50.^{m}9 + 3^{o}58′ C/1999 G3
3.496 0 56.3 + 4 31  C/1999 G4
28.729 2 27.6 +12 59  C/1999 H6

NOVA IN THE LARGE MAGELLANIC CLOUD 2000 A. Gilmore reports the following photometry obtained by P. M. Kil-

martin with the 0.6-m f/16 reflector at Mount John Observatory at Lake Tekapo (reference star CPD −70^{o}399, calibrated by Cousins): July 16.29 UT, V = 12.83 ± 0.02, U − B = −0.91 ± 0.08, B − V = −0.12 ± 0.00, V − R = +1.15 ± 0.05, V − I = +0.77 ± 0.03, airmass 1.97.

2000 July 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green